# CursorCore: Assist Programming through Aligning Anything

### v2

@@ Abstract
Large language models have been successfully applied to programming assistance tasks, such as code completion, code insertion, and instructional code editing. However, these applications remain insufficiently automated and struggle to effectively integrate various types of information during the programming process, including coding history, current code, and user instructions. In this work, we propose a new conversational framework that comprehensively integrates these information sources, collect data to train our models and evaluate their performance. Firstly, to thoroughly evaluate how well models align with different types of information and the quality of their outputs, we introduce a new benchmark, APEval (Assist Programming Eval), to comprehensively assess the performance of models in programming assistance tasks. Then, for data collection, we develop a data generation pipeline, Programming-Instruct, which synthesizes training data from diverse sources, such as GitHub and online judge platforms. This pipeline can automatically generate various types of messages throughout the programming process. Finally, using this pipeline, we generate 219K samples, fine-tune multiple models, and develop the CursorCore series. We show that CursorCore outperforms other models of comparable size. This framework unifies applications such as inline chat and automated editing, contributes to the advancement of coding assistants. Code, models and data are freely available at https://github.com/TechxGenus/CursorCore.
@@ 1 Introduction
Since the rise of large language models (LLMs), AI-assisted programming technology has developed rapidly, with many powerful LLMs being applied in this field Zan et al. (2022); Liang et al. (2024); Yang et al. (2024). The technology mainly takes two forms. One form involves completing a specified code snippet at the end or inserting corresponding code at a designated position, typically accomplished by foundation models Chen et al. (2021); Bavarian et al. (2022) that support relevant input formats. The other form involves generating or editing code snippets based on natural language instructions or reflections through interaction with the environment, usually carried out by instruction models that have been further aligned Shinn et al. (2023); Cassano et al. (2023b); Muennighoff et al. (2024); Paul-Gauthier (2024). Figure 1 shows simple examples of these forms.
@@
To solve these issues, in this paper, we introduce a new framework of AI-assisted programming task: Assistant-Conversation to align anything during programming process. To comprehensively evaluate the alignment of models with different information in the programming process and the quality of the corresponding outputs, we propose a new benchmark, APEval (Assist Programming Eval), to comprehensively assess the performance of models in assisting programming. For the Assistant-Conversation framework, we build a data generation pipeline, Programming-Instruct, to synthesize corresponding training data from various data sources. This data generation method can produce any types of messages throughout the programming process, without any additional human annotation and does not rely on specific models. We use this pipeline to generate 219K data points and use them to fine-tune multiple models, resulting in the CursorCore series. These models achieve state-of-the-art results when compared with other models of comparable size.
@@
  • APEval: A benchmark for assessing the ability to utilize various types of information to assist programming.
@@ 2 Assistant-Conversation: New Conversation Framework for Programming Assistants
In this section, we introduce a new conversational framework, Assistant-Conversation, aimed at simplifying the programming process. The framework leverages all available information during programming to streamline work for programmers. By precisely defining various types of information and their formats, Assistant-Conversation directly aligns with the input and output requirements of applications such as automated editing and inline chat. This framework facilitates model alignment, enabling fast and accurate generation and parsing.
@@ 2.1 Framework Formulation
We introduce the constituent elements of Assistant-Conversation: System (S), History (H), Current (C), User (U), and Assistant (A). The Assistant (A) represents the output of the model, while the inputs consist of the System (S), History (H), Current (C), and User (U). Figures 1 and 2 shows several examples of them. These definitions will be referenced throughout the rest of this work.
@@
Current C
The code currently being processed, along with temporary information like cursor position or selected code area.
User U (Optional)
@@
Figure 2: Examples of Assistant-Conversation from our training data. The top example demonstrates predicting the corresponding edits and explanations based on historical edits and the current code. The bottom example demonstrates predictions based on the current code and user instructions.
@@ 2.3 Specifications and Implementation
To represent a piece of code like C, we can either use it directly or wrap it in a markdown code block. However, representing code changes, such as H or changes in A, is more complex. We can either use the whole code, patches that alter the code, or records of both the modification locations and the specific changes. Some methods work well but experience issues when handling longer texts, such as outputting the entire modified code, which can be slow. Other methods output minimal content, like providing only the modification locations and changes. These are faster but still not optimal in terms of performance. We represent code changes in the experiments of the main body using the whole code format, and we investigate different ways to represent these modifications, as detailed in Appendix B. Additionally, we explore methods for compressing historical code changes in Appendix G.
In some cases, programmers assign assistants to focus on specific areas of code. They might use the cursor to mark a general location or directly select a range of code, as shown in Figure 2. We handle this by treating them as special tokens (see Appendix E for further details).
We structure conversations in the order of S-H-C-U-A to match the actual workflow. This mirrors the chronological sequence in which information is generated during the programming process. By doing so, we maximize prefix overlap across multiple requests, utilizing prefix caching to reduce redundant kv-cache computations and improve efficiency Zheng et al. (2023a). A is organized in code-chat order, prioritizing code edits due to their importance in real-time applications where speed is crucial.
@@ 3.1 Benchmark overview
Table 1: APEval Statistics. Present statistical information about H, C, and U in our benchmark.
Statistics Sample Num
Total 164
Each type 41 / 41 / 41 / 41
H Statistics Mean / Max
Num (Snippets) 2.8 / 10
Num (Lines) 21.7 / 139
Num (Chars) 0.6K / 5.1K
C Statistics Mean / Max
Num (Lines) 8.4 / 31
Num (Chars) 0.3K / 1.4K
U Statistics Mean / Max
Num (Lines) 3.2 / 19
Num (Chars) 0.2K / 1.2K
Past benchmarks assessing LLM code capabilities have effectively evaluated tasks like program synthesis Chen et al. (2021); Austin et al. (2021), code repair Muennighoff et al. (2024); Jimenez et al. (2024), and instructional code editing Cassano et al. (2023b); Paul-Gauthier (2024); Guo et al. (2024b). However, they fall short in fully assessing how models use various types of information to assist in programming. This gap calls for a new benchmark.
As discussed in Section 2.1, programming assistance can involve different types of information, with H and U being optional. Thus, there are four possible combinations of information: H, C, U; H, C; C, U; and only C. HumanEval Chen et al. (2021) is a well-known benchmark for evaluating code completion. It has been extended to assess other tasks such as code insertion Bavarian et al. (2022), instruction-based tasks CodeParrot (2023); Muennighoff et al. (2024), and multilingual generation Zheng et al. (2023b); Cassano et al. (2023a). We refer to these works and further extend it to comprehensively evaluate the model’s ability to assist programming. We randomly categorize each task into one of the four types, then manually implement the functions and simulate the potential instructions that programmers might give to an LLM during the process, collecting all interactions. We invite programmers with varying levels of experience to annotate the data. After processing, we get the new benchmark, Assist Programming Eval (APEval). Detailed statistics are shown in Table 1. Specific details regarding the collection process and examples of our benchmark can be found in Appendix C.
@@ 3.2 Evaluation Process and Metrics
In all tasks, we use the classic Pass@1 metric to execute the generated code, which is the simplest version of the Pass@k metric Chen et al. (2021). Since APEval is an extension of HumanEval, we use the test set created by EvalPlus Liu et al. (2023). We report the results from both the basic and extra tests. We provide the model with relevant information during the programming process, and the model immediately returns the modified code. Some methods may improve performance by increasing the number of output tokens to model the thinking process; we discuss this further in Appendix F.
@@ 4 Programming-Instruct: Collect any data during programming
To align models with programming-related data, relevant training data must be collected. While large amounts of unsupervised code Kocetkov et al. (2023) and instruction data Wei et al. (2023b); Luo et al. (2024b) have been gathered, there remains a significant lack of data on the coding process. Manually annotating the coding process is expensive, so we propose Programming-Instruct, a method to automate this data collection.
@@ 4.1 Data Sources
To ensure both quality and diversity in the coding process data, we collect information from three different sources: AIprogrammer, Git commit, and Online Submit.
@@
For each code snippet, we use LLMs to generate the corresponding coding history. Since human coding approaches vary widely, we utilize several LLMs, each guided by three distinct prompts, representing novice, intermediate, and expert programmers. The LLMs then return their version of the coding process. Prompts used are shown in Appendix L.
@@
Online Submit
Many online coding platforms like Leetcode allow users to submit code for execution and receive feedback. During this process, users continuously modify their code until it is finalized. We also make use of this data.
@@
Figure 3: Samples from AIprogrammer, Git Commit and Online Submit.
@@ 4.2 Data Processing
After collecting a large number of coding processes, we process them to meet the requirements of Assistant-Conversation. Figure 4 shows the steps of data processing. First, we randomly select a time point in the coding process, referred to as C. As mentioned in Section 2.1, H and U are optional, we need to collect four types of data distinguished according to input data types: H, C, U; H, C; C, U; and only C. For each sample, we randomly designate one type. If the selected type includes H, We use the preceding edits of C as the historical records H.
We then handle each type of data based on whether U is available. For cases without U, we segment the changes from C to F based on continuity, referring to them as M, and let LLMs judge whether each segment of M aligns with user’s purpose through principle-driven approaches Bai et al. (2022); Sun et al. (2023); Lin et al. (2024). This approach accounts for ambiguity in user intent when inferring from H or C. For example, if a programmer actively adds some private information at the beginning of the code without it being mentioned in the previous records, LLMs should not predict this change. We discard segments deemed irrelevant, and merge the remaining ones as outputs that models need to learn to predict. For cases with U, we follow the instruction generation series methods Wang et al. (2023); Wei et al. (2023b); Luo et al. (2024b) by inputting both the historical edits and current code into the LLM, prompting it to generate corresponding instructions.
@@
Figure 4: Data processing pipeline. The randomly selected time point is the third, and the randomly selected data type is H and C.
5 CursorCore: Fine-tune LLMs to align anything
5.1 Base models
We fine-tune existing base LLMs to assist with programming tasks. Over the past few years, many open-source foundation models have been trained on large code corpora sourced from GitHub and other platforms, demonstrating strong performance in coding. We choose the base versions of Deepseek-Coder Guo et al. (2024a), Yi-Coder AI et al. (2024) and Qwen2.5-Coder Hui et al. (2024) series, as fine-tuning is generally more effective when applied to base models rather than instruction models. After training, we refer to them as CursorCore-DS, CursorCore-Yi and CursorCore-QW2.5 series. Deepseek-Coder has achieved state-of-the-art performance on numerous coding-related benchmarks over the past year, gaining wide recognition. Yi-Coder and Qwen2.5-Coder are the most recently released models at the start of our experiments and show the best performance on many benchmarks for code now. These models are widely supported by the community, offering a good balance between size and performance, making them suitable for efficient experimentation. For ablation experiments, we use the smallest version, Deepseek-Coder-1.3B, to accelerate the process. We use a chat template adapted from ChatML OpenAI (2023) to model Assistant-Conversation during training, as detailed in Appendix J.
5.2 Training data
@@
Online Submit 60.5K 44 3.8 / 96 4.8K / 357.2K 1.9K / 35.1K
We use Programming-Instruct to collect data. For AIprogrammer, we gather code snippets from datasets such as the stack Kocetkov et al. (2023) and oss-instruct Wei et al. (2023b), then prompt LLMs to generate the programming process. For Git commit data, we collect relevant information from editpackft Cassano et al. (2023b) (a filtered version of commitpackft Muennighoff et al. (2024)) and further refine it through post-processing and filtering. Regarding online submission data, we source the programming process from the Codenet dataset Puri et al. (2021). First, we group all submissions by user for each problem, then exclude invalid groups without correct submissions to obtain complete programming processes. These are then fed into the processing pipeline to generate the final training data. In total, we accumulate 219K samples, with detailed statistics and distributions shown in Tables 2, 3, 6, 6, 8 and 8. AIprogrammer data has the shortest average length, while Online Submit data has the longest. To ensure compatibility with previous chatbot-style interactions and further improve model performance, we also incorporate the evol-instruct dataset ISE-UIUC (2023) collected using the GPT series Ouyang et al. (2022), which has been widely recognized for its high quality during training. Following StarCoder’s data processing approach Li et al. (2023), we decontaminate our training data.

### v3

@@ Abstract
Large language models have been successfully applied to programming assistance tasks, such as code completion, code insertion, and instructional code editing. However, these applications remain insufficiently automated and struggle to effectively integrate various types of information during the programming process, including coding history, code context, and user instructions. In this work, we propose a new framework that comprehensively integrates these information sources, and collect data to train models and evaluate their performance. Firstly, to thoroughly evaluate how well models align with different types of information and the quality of their outputs, we introduce a new benchmark, APEval (Assist Programming Eval), to comprehensively assess the performance of models in programming assistance tasks. Then, for data collection, we develop a data generation pipeline, Programming-Instruct, which synthesizes training data from diverse sources, such as GitHub and online judge platforms. This pipeline can automatically generate various types of messages throughout the programming process. Finally, using this pipeline, we generate 219K samples, fine-tune multiple models, and develop the CursorCore series. We show that CursorCore outperforms other models of comparable size. This framework unifies applications such as inline chat and automated editing, contributes to the advancement of coding assistants.
Large Language Models for Code, AI-Assisted Programming
@@ 1 Introduction
Since the rise of large language models (LLMs), AI-assisted programming technology has developed rapidly, with many powerful LLMs being applied in this field (Zan et al., 2022; Liang et al., 2024; Yang et al., 2024). The technology mainly takes two forms. One form involves completing a specified code snippet at the end or inserting corresponding code at a designated position, typically accomplished by foundation models (Chen et al., 2021; Bavarian et al., 2022) that support relevant input formats. The other form involves generating or editing code snippets based on natural language instructions or reflections through interaction with the environment, usually carried out by instruction models that have been further aligned (Shinn et al., 2023; Cassano et al., 2023b; Muennighoff et al., 2024; Paul-Gauthier, 2024). Figure 1 shows simple examples of these forms.
@@
To solve these issues, in this paper, we introduce a new framework of AI-assisted programming task: Assistant-Conversation to align anything during programming process. To comprehensively evaluate the alignment of models with different information in the programming process and the quality of the corresponding outputs, we propose a new benchmark, APEval (Assist Programming Eval), to comprehensively assess the performance of models in assisting programming. For the Assistant-Conversation framework, we build a data generation pipeline, Programming-Instruct, to synthesize corresponding training data from various data sources. This data generation method can produce any types of messages throughout the programming process, without any additional human annotation and does not rely on specific models. We use it to generate 219K data points and use them to fine-tune multiple models, resulting in the CursorCore series. These models achieve state-of-the-art results when compared with other models of comparable size.
@@
  • APEval: A comprehensive benchmark for assessing the ability to utilize various types of information to assist programming.
@@
Code, models and data are freely available at https://github.com/TechxGenus/CursorCore.
@@ 2 Assistant-Conversation: New Conversation Framework for Programming Assistants
In this section, we introduce a new conversational framework, Assistant-Conversation, aimed at simplifying the programming process [^1] In this work, “conversation” refers to the common format used in LLM generation, rather than multi-turn dialogues.. The framework leverages all available information during programming to streamline work for programmers. By precisely defining various types of information and their formats, Assistant-Conversation directly aligns with the input and output requirements of applications such as automated editing and inline chat. This framework facilitates model alignment, enabling fast and accurate generation and parsing.
@@ 2.1 Framework Formulation
Figure 2: Examples of Assistant-Conversation from our training data. The top example demonstrates predicting the corresponding edits and explanations based on historical edits and the current code context. The bottom example demonstrates predictions based on the current code and user instructions.
We introduce the elements of Assistant-Conversation: System (S), History (H), Current Context (C), User Instruction (U), and Assistant Output (A). A represents the output of the model, while the inputs consist of S, H, C, U. Figures 1 and 2 shows several examples of them. These definitions will be referenced throughout the rest of this work.
@@
Current Context C
The code context currently being processed, along with temporary information like cursor position or selected code area.
User Instruction U (Optional)
@@ 2.3 Specifications and Implementation
To represent a piece of code like C, we can either use it directly or wrap it in a markdown code block. However, representing code changes, such as H or changes in A, is more complex. We can either use the whole code, patches that alter the code, or records of both the modification locations and the specific changes. Some methods work well but experience issues when handling longer texts, such as outputting the entire modified code, which can be slow. Other methods output minimal content, like providing only the modification locations and changes. These are faster but still not optimal in terms of performance. We represent code changes in the experiments of the main body using the whole code format, and we investigate different ways to represent these modifications, as detailed in Appendix B. Additionally, we explore methods for compressing historical code changes in Appendix I.
In some cases, programmers assign assistants to focus on specific areas of code. They might use the cursor to mark a general location or directly select a range of code, as shown in Figure 2. We handle this by treating them as special tokens (see Appendix F for further details).
We structure conversations in the order of S-H-C-U-A to match the actual workflow. This mirrors the chronological sequence in which information is generated during the programming process. By doing so, we maximize prefix overlap across multiple requests, utilizing prefix caching to reduce redundant kv-cache computations and improve efficiency (Zheng et al., 2023a). A is organized in code-chat order, prioritizing code edits due to their importance in real-time applications where speed is crucial.
@@ 3.1 Benchmark overview
Past benchmarks assessing LLM code capabilities have effectively evaluated tasks like program synthesis (Chen et al., 2021; Austin et al., 2021), code repair (Muennighoff et al., 2024; Jimenez et al., 2024), and instructional code editing (Cassano et al., 2023b; Paul-Gauthier, 2024; Guo et al., 2024b). However, they fall short in fully assessing how models use various types of information to assist in programming. This gap calls for a new benchmark.
Table 1: APEval Statistics and breakdown of tasks by information type.
APEval Statistics
Python 164 Samples
Multilingual 984 Samples
Language Python, C++, Java, JavaScript, Go, Rust
Details Mean Max
Snippets (H) 2.8 10
[…]
As discussed in Section 2.1, programming assistance can involve different types of information, with H and U being optional. Thus, there are four possible combinations of information: H, C, U; H, C; C, U; and only C. HumanEval (Chen et al., 2021) is a well-known benchmark for evaluating code completion. It has been extended to assess other tasks such as code insertion (Bavarian et al., 2022), instruction-based tasks (CodeParrot, 2023; Muennighoff et al., 2024), and multilingual generation (Zheng et al., 2023b; Cassano et al., 2023a). We refer to these works and further extend it to comprehensively evaluate the model’s ability to assist programming. We randomly categorize each task into one of the four types, then manually implement the functions and simulate the potential instructions that programmers might give to an LLM during the process, collecting all interactions. We invite programmers with varying levels of experience to annotate the data. After processing, we get the new benchmark, Assist Programming Eval (APEval), which contains approximately 1K multilingual samples. Detailed statistics are shown in Table 1. Specific details regarding the collection process and examples of our benchmark can be found in Appendix C, which includes detailed human annotation rubric and results.
@@ 3.2 Evaluation Process and Metrics
In all tasks, we use the classic Pass@1 metric to execute the generated code, which is the simplest version of the Pass@k metric (Chen et al., 2021). Since APEval is an extension of HumanEval, we evaluate its Python version using the test set created by EvalPlus (Liu et al., 2023) and assess its other language versions using bigcode-evaluation-harness (Ben Allal et al., 2022). We set the Python version as the default version for evaluation, and report the results from both the basic and extra tests. We provide the model with relevant information during the programming process, and the model immediately returns the modified code. Some methods may improve performance by increasing the number of output tokens to model the thinking process; we discuss this further in Appendix G.
@@ 4 Programming-Instruct: Collect any data during programming
To align models with programming-related data, relevant training data must be collected. While large amounts of unsupervised code (Kocetkov et al., 2023) and instruction data (Wei et al., 2023b; Luo et al., 2024b) have been gathered, there remains a significant lack of data on the coding process. Manually annotating the coding process is expensive, so we propose Programming-Instruct, a method to automate this data collection.
@@ 4.1 Data Sources
Figure 3: Samples from AI Programmer, Git Commit and Online Judge Submission.
Figure 4: Data processing pipeline. The randomly selected time point is the third, data type is H and C.
To ensure both quality and diversity in the coding process data, we collect information from three different sources: AI Programmer, Git Commit, and Online Judge submission.
@@
For each code snippet, we use LLMs to generate the corresponding coding history. Since human coding approaches vary widely, we utilize several LLMs, each guided by three distinct prompts, representing novice, intermediate, and expert programmers. The LLMs then return their version of the coding process. Prompts used are shown in Appendix O.
@@
Online Judge Submission
Many online coding platforms like Leetcode and Codeforces allow users to submit code for execution and receive feedback. During this process, users continuously modify their code until it is finalized. We also make use of this data.
@@ 4.2 Data Processing
After collecting programming processes, we process them to meet the requirements of Assistant-Conversation. Figure 4 shows the steps of data processing. First, we randomly select a time point in the coding process, referred to as C. As mentioned in Section 2.1, H and U are optional, we need to collect four types of data distinguished according to input data types: H, C, U; H, C; C, U; and only C. For each sample, we randomly designate one type. If the selected type includes H, We use the preceding edits of C as the historical records H.
We then handle each type of data based on whether U is available. For cases without U, we segment the changes from C to F based on continuity, referring to them as M, and let LLMs analyze and then judge whether each segment of M aligns with user’s purpose through principle-driven approaches (Bai et al., 2022; Sun et al., 2023; Lin et al., 2024). This approach accounts for ambiguity in user intent when inferring from H or C. For example, if a programmer actively adds some private information at the beginning of the code without it being mentioned in the previous records, LLMs should not predict this change. We discard segments deemed irrelevant, and merge the remaining ones as outputs that models need to learn to predict. For cases with U, we follow the instruction generation series methods (Wang et al., 2023b; Wei et al., 2023b; Luo et al., 2024b) by inputting both the historical edits and current code into the LLM, prompting it to generate corresponding instructions.
@@
Online Judge Submission 60.5K 44 3.8 / 96 4.8K / 357.2K 1.9K / 35.1K
5 CursorCore: Fine-tune LLMs to align anything
5.1 Base models
We fine-tune existing base LLMs to assist with programming tasks. Over the past few years, many open-source foundation models have been trained on large code corpora sourced from GitHub and other platforms, demonstrating strong performance in coding. We choose the base versions of Deepseek-Coder (Guo et al., 2024a), Yi-Coder (AI et al., 2024) and Qwen2.5-Coder (Hui et al., 2024) series, as fine-tuning is generally more effective when applied to base models rather than instruction models. After training, we refer to them as CursorCore-DS, CursorCore-Yi and CursorCore-QW2.5 series. Deepseek-Coder has achieved state-of-the-art performance on numerous coding-related benchmarks over the past year, gaining wide recognition. Yi-Coder and Qwen2.5-Coder are the most recently released models at the start of our experiments and show the best performance on many benchmarks for code now. These models are widely supported by the community, offering a good balance between size and performance, making them suitable for efficient experimentation. For ablation experiments, we use the smallest version, Deepseek-Coder-1.3B, to accelerate the process. We use a chat template adapted from ChatML (OpenAI, 2023) to model Assistant-Conversation during training, as detailed in Appendix M. Training details can be found in Appendix E.
Figure 5: Distribution of programming language in the training data.
Figure 6: Distribution of history snippet counts in the training data.
5.2 Training data
Figure 7: Distribution of input lengths in the training data.
Figure 8: Distribution of output lengths in the training data.
We use Programming-Instruct to collect data. For AI Programmer, we gather code snippets from datasets such as the Stack (Kocetkov et al., 2023) and OSS-Instruct (Wei et al., 2023b), then prompt LLMs to generate the programming process. For Git Commit data, we collect relevant information from EditPackFT (Cassano et al., 2023b) (a filtered version of CommitPackFT (Muennighoff et al., 2024)) and further refine it through post-processing and filtering. Regarding Online Judge Submission data, we source the programming process from the Codenet dataset (Puri et al., 2021). First, we group all submissions by user for each problem, then exclude invalid groups without correct submissions to obtain complete programming processes. These are then fed into the processing pipeline to generate the final training data. In total, we accumulate 219K samples, with detailed statistics and distributions shown in Tables 2, 3, 5, 6, 7 and 8. AI Programmer data has the shortest average length, while Online Judge Submission data has the longest. To ensure compatibility with previous chatbot-style interactions and further improve model performance, we also incorporate the Evol-Instruct dataset (ISE-UIUC, 2023) collected using the GPT series (Ouyang et al., 2022), which has been widely recognized for its high quality during training. Following StarCoder’s data processing approach (Li et al., 2023), we decontaminate our training data.
During data collection, we randomly utilize two powerful open-source LLMs: Mistral-Large-Instruct (Mistral-AI, 2024b) and Deepseek-Coder-V2-Instruct (DeepSeek-AI et al., 2024). These models have demonstrated performance comparable to strong closed-source models like GPT-4o across many tasks, and are currently the only two open-source models scoring over 90% on the classic HumanEval benchmark at the start of our experiment. Additionally, they are more cost-effective and offer easier reproducibility than GPT-4o. For Mistral-Large-Instruct, we quantize the model using the GPTQ (Frantar et al., 2022) algorithm and deploy it locally with SGLang (Zheng et al., 2023a) and Marlin kernel (Frantar et al., 2024) on 4 Nvidia RTX 4090 GPUs. For Deepseek-Coder-V2-Instruct, we use its official API for integration.